# Efficient Computation of Voronoi Diagrams Using Point-in-Cell Tests

### v1

@@ 3.2. Edge/face flags of polygon/polytope
Our task is to compute a set of triangle/tetrahedron-cell intersections as fast as possible. This requires an efficient search for pairs of triangles/tetrahedra and Voronoi cells that intersect with each other. Such triangle/tetrahedron-cell pairs can be easily found by a propagation approach [Yan et al., 2009, 2013].
@@
The intermediate and final results of each triangle/tetrahedron-cell intersection are both a polygon/polytope, and we can attach a flag to each edge/face of the polygon/polytope to achieve the propagation. For an edge/face of polygon/polytope $\mathcal{P}$, if it is shared by $\mathcal{P}$ and a domain triangle/tetrahedron $\mathcal{T}_{b}$, its flag is set as $(-1)\times$ the index of $\mathcal{T}_{b}$, i.e., b-b, while if it is shared by $\mathcal{P}$ and a Voronoi cell $\Omega_{j}$, its flag is set as the index of the cell, i.e., $j$. Thus, the edge/face flags of each triangle/tetrahedron $\mathcal{T}_{a}\in\mathcal{M}$ can be initialized by using the indices of the adjacent triangles/tetrahedra, see Fig.2(a). If $\mathcal{T}_{a}$ intersects with the Voronoi cell $\Omega_{i}$, the intersection $\mathcal{P}_{a,i}$ can be generated after several clippings. Each valid clipping will produce a new edge/face, its flag is set as the index of the neighboring site related to the clipping, see Fig.2(b). Once the intersection $\mathcal{P}_{a,i}$ is finally generated, its edge/face flags indicate either the neighboring Voronoi cells of $\Omega_{i}$ intersected with $\mathcal{T}_{a}$, or the adjacent triangles/tetrahedra of $\mathcal{T}_{a}$ intersected with $\Omega_{i}$, so that other triangle/tetrahedron-cell pairs can be obtained quickly using edge/face flags of $\mathcal{P}_{a,i}$.
(a) edge/face flags of $\mathcal{T}_{a}\in\mathcal{M}$
(b) edge/face flags after clippings
Figure 2. Each edge/face flag indicates the adjacent domain triangle/tetrahedron (<0<0) or the neighboring Voronoi cell (>0>0).
@@
Figure 3. Illustration of the edge-based bisector search. Initially, $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\in\Omega_{j}$ (a), we iteratively compute the intersecting point $\mathbf{q}$ (yellow) of $\mathbf{p}_{s}\mathbf{p}_{t}$ (pink) and the bisector $B_{i,j}$ (purple), and update $j$ as the index of the nearest site of $\mathbf{q}$ (b-c), until $\mathbf{q}\in\Omega_{i}$, then the target bisector is found (d). Using the target bisector to clip the polygon/polytope, the two endpoints of the rest of the edge are both in $\Omega_{i}$ (e). Each blue arrow shows the nearest site $\mathbf{v}$ of a point $\mathbf{p}$. $i=1$ in this figure.
4. The Proposed Method
Benefiting from the convexity of Voronoi cells, it is known that a polygon/polytope lies entirely in a Voronoi cell if all its corner points are in the cell. If a triangle/tetrahedron $\mathcal{T}_{a}\in\mathcal{M}$ intersects with a Voronoi cell $\Omega_{i}$, our strategy is to remove the part of $\mathcal{T}_{a}$ outside $\Omega_{i}$, ensuring all the corner points of the rest are in $\Omega_{i}$.
Each triangle/tetrahedron can be decomposed into smaller polygons/polytopes, each belonging to a Voronoi cell, the details are provided in 4.4. Because all Voronoi cells are unknown, how to judge whether a space point is in a target Voronoi cell is the key to our method. We set up a test called point-in-cell in 4.1 to accomplish the judgment. By using the tests, we present a novel solution that only involves valid clippings to generate triangle/tetrahedron-cell intersection in 4.3, and the bisector of each clipping can be found by an edge-based search scheme introduced in 4.2.
4.1. The point-in-cell test
Given a point $\mathbf{p}$ and the index $i$ of a Voronoi cell $\Omega_{i}$, the point-in-cell test tells us whether the point is in the cell or not, which is frequently invoked in our method.
Based on the definition of Voronoi cells, $\mathbf{p}$ is in $\Omega_{i}$ if the distance from $\mathbf{p}$ to $\mathbf{v}_{i}$ is not greater than that from $\mathbf{p}$ to any other sites. Let $\mathcal{N}_{\mathbf{p}}$ be the nearest site of $\mathbf{p}$. We can conclude that $\mathbf{p}\in\Omega_{i}$ if $\mathcal{N}_{\mathbf{p}}=\mathbf{v}_{i}$ or if the following equation holds:
(2) $\|\mathbf{p}-\mathbf{v}_{i}\|=\|\mathbf{p}-\mathcal{N}_{\mathbf{p}}\|.$
Thus, only one time of nearest neighbor search is required in each point-in-cell test. Owing to the accuracy issue of floating numbers in computers, we set a tolerance $\xi$ to evaluate equation (2), i.e.,
(3) $\|\mathbf{p}-\mathbf{v}_{i}\|-\|\mathbf{p}-\mathcal{N}_{\mathbf{p}}\|<\xi,$
where $\xi$ is a small positive number, like $10^{-8}$, and can be adjusted with the scale of applications.
4.2. Edge-based bisector search
It is known that the points shared by two neighboring Voronoi cells $\Omega_{u},\Omega_{w}$ must be on the bisector $B_{u,w}$ of the two corresponding sites $\mathbf{v}_{u},\mathbf{v}_{w}$, and the clipping using $B_{u,w}$ must contribute to the generation of Voronoi cells $\Omega_{u},\Omega_{w}$. This observation inspires us the way to search for the target bisector on edge $\mathbf{p}_{s}\mathbf{p}_{t}$.
Initially, $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\in\Omega_{j}$, we construct the bisector $B_{i,j}$ of the corresponding sites $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, and compute the intersecting point $\mathbf{q}$ of $\mathbf{p}_{s}\mathbf{p}_{t}$ and $B_{i,j}$. By using the point-in-cell test, we can easily obtain the relationship between $\mathbf{q}$ and $\Omega_{i}$: (1) If the test answers false, and the nearest site of $\mathbf{q}$ is denoted as $\mathbf{v}_{k}$, then $k$ must be different with $i$ and $j$, which means that the clipping using $B_{i,j}$ must be useless. In this case, we update $j\leftarrow k,\mathbf{p}_{t}\leftarrow\mathbf{q}$ and repeat the above steps, i.e., update $B_{i,j}$ and $\mathbf{q}$, and redo the point-in-cell test. (2) If the test answers true, we know that the whole segment $\mathbf{p}_{s}\mathbf{q}$ is in the cell $\Omega_{i}$, then the newest $B_{i,j}$ is the target bisector, which can be directly used to clip the polygon/polytope. Fig. 3 shows an illustration.
The target bisector on the edge can be found after several point-in-cell tests. The number of tests is proportional to the number of Voronoi cells intersecting the edge. To reduce the number of tests, a minor modification can be made before searching for the target bisector, which will be discussed later in 6.1.
@@ 4.2. Edge-based bisector search
Figure 4. Illustration of generating the intersection $\mathcal{P}$ of a triangle/tetrahedra $\mathcal{T}\in\mathcal{M}$ and the Voronoi cell $\Omega_{i}$. At least one corner point of the polygon/polytope should be in $\Omega_{i}$. Select an edge whose starting point is in $\Omega_{i}$ but the end point is not, then find a bisector by using the proposed edge-based search scheme to clip the polygon/polytope. Repeat the selection and clipping until all the corner points are in $\Omega_{i}$. $i=1$ in this figure.
@@
Figure 5. Illustration of generating the intersection $\mathcal{P}_{a,j}$ of a triangle/tetrahedron $\mathcal{T}_{a}\in\mathcal{M}$ and the Voronoi cell $\Omega_{j}$, where there is no corner point of $\mathcal{T}_{a}$ being in $\Omega_{j}$. (a) The intersection $\mathcal{P}_{a,u}$ (light blue polygon) of $\mathcal{T}_{a}$ and $\Omega_{u}$ has been obtained, one of its edges/faces (red segment) involves a clipping related to the bisector $B_{u,j}$, and its adjacent edges/faces (blue segments) are useful to construct an intermediate polygon/polytope $\mathcal{G}_{a,j}$; (b-c) construction of $\mathcal{G}_{a,j}$; (d) generation of $\mathcal{P}_{a,j}$ from $\mathcal{G}_{a,j}$ by using Algorithm 1. $u=1,j=2$ in this figure.
0: a triangle/tetrahedron $\mathcal{T}\in\mathcal{M}$ and a site index ii.
@@
3:  select an edge $\mathbf{p}_{s}\mathbf{p}_{t}$ from $\mathcal{P}$, requiring that $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\notin\Omega_{i}$;
4:  find a bisector $B_{i,j}$ on $\mathbf{p}_{s}\mathbf{p}_{t}$ (section 4.2);
@@
ALGORITHM 1 Generation of triangle/tetrahedron-cell intersection
4.3. Generation of triangle/tetrahedron-cell intersection
Suppose that at least one corner point of a triangle/tetrahedron $\mathcal{T}$ is in the Voronoi cell $\Omega_{i}$ of site $\mathbf{v}_{i}$, we present a novel solution for generating the intersection $\mathcal{P}$ of $\mathcal{T}$ and $\Omega_{i}$. Specifically, $\mathcal{P}$ is initialized as $\mathcal{T}$, and it is iteratively removed the part outside $\Omega_{i}$. The steps are as follows: (1) Select an edge $\mathbf{p}_{s}\mathbf{p}_{t}$ from $\mathcal{P}$, requiring that $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\notin\Omega_{i}$. (2) Find a bisector $B_{i,j}$ by using the proposed edge-based search scheme on $\mathbf{p}_{s}\mathbf{p}_{t}$, and use it to clip $\mathcal{P}$, the result is still denoted as $\mathcal{P}$. Repeat the selection and clipping steps until $\mathcal{P}$ has no edge meeting the requirement in step 1, which implies that all the corner points of $\mathcal{P}$ are in $\Omega_{i}$, i.e., the triangle/tetrahedron-cell intersection is finally obtained. The steps are also listed in Algorithm 1, and an illustration is given in Fig. 4.
At most one clipping is involved for each edge of the intersection, and all the clippings are contributed to the final result. This is a great advantage of our solution, since a part of the clippings in both DT-based and kNN-based methods could be useless for generating such intersections.
4.4. Computation of clipped or restricted Voronoi diagram
We now concentrate on the generation of clipped or restricted Voronoi cells $\left\{\Omega_{i|\mathcal{M}}\right\}_{i=1}^{n}$. Initially, they are set to empty, i.e., $\Omega_{i|\mathcal{M}}=\emptyset,i=1,\dots,n$. We can decompose each domain triangle/tetrahedron into smaller polygons/polytopes, each of which is entirely in a Voronoi cell such that it can be added to the corresponding clipped or restricted cell. Since the decomposition process is identical for all triangles/tetrahedra, we take $\mathcal{T}_{a}\in\mathcal{M}$ as an example to describe our method. The pseudo-code can be found in Algorithm 2.
4.4.1. First polygon/polytope
4.4.2. Propagation
It has to further compute the intersections of $\mathcal{T}_{a}$ and other Voronoi cells, the triangle/tetrahedron-cell pairs can be obtained by using the edge/face flags of $\mathcal{P}_{a,i}$. Specifically, We set up a queue $Q$ to store the unvisited sites whose Voronoi cells are intersected with $\mathcal{T}_{a}$. After the first polygon/polytope $\mathcal{P}_{a,i}$ is generated, we traverse its all edges/faces. For each edge/face, if it has a positive flag $k$, meaning that $\Omega_{k}$ must intersect with $\mathcal{T}_{a}$, we add $\mathbf{v}_{k}$ to the queue. Iteratively, a site $\mathbf{v}_{j}$ is popped from $Q$, and we compute the intersection $\mathcal{P}_{a,j}$ of $\mathcal{T}_{a}$ and $\Omega_{j}$, which is then added to $\Omega_{j|\mathcal{M}}$. The queue $Q$ can be further updated by adding more unvisited sites using the edge/face flags of $\mathcal{P}_{a,j}$. The procedure is terminated once the queue is empty, and the decomposition on $\mathcal{T}_{a}$ is completed.
@@ 4.4.2. Propagation
1: build a structure $\mathcal{S}$ of $\mathcal{V}$ for nearest neighbor search (section 6.2);
2: find the nearest site for each corner point of the domain;
@@
4:for each $\mathcal{T}_{a}\in\mathcal{M}$ do
@@
5:  $\mathbf{v}_{i}\leftarrow$ the nearest site of the first corner point of $\mathcal{T}_{a}$;
6:  compute the intersection $\mathcal{P}_{a,i}$ of $\mathcal{T}_{a}$ and $\Omega_{i}$ using Algorithm 1;
7:  $\Omega_{i|\mathcal{M}}\leftarrow\Omega_{i|\mathcal{M}}\bigcup\mathcal{P}_{a,i}$;
8:  $Q\leftarrow\emptyset$;
9:  add other sites to $Q$ using edge/face flags of $\mathcal{P}_{a,i}$;
10:  while $Q$ is not empty do
11:   $\mathbf{v}_{j}\leftarrow$ popped from $Q$;
12:   construct an intermediate polygon/polytope $\mathcal{G}_{a,j}$ from $\mathcal{T}_{a}$;
13:   compute the intersection $\mathcal{P}_{a,j}$ of $\mathcal{G}_{a,j}$ and $\Omega_{j}$ using Algorithm 1;
14:   $\Omega_{j|\mathcal{M}}\leftarrow\Omega_{j|\mathcal{M}}\bigcup\mathcal{P}_{a,j}$;
15:   add other unvisited sites to $Q$ using edge/face flags of $\mathcal{P}_{a,j}$;
16:  end while
17:end for
ALGORITHM 2 Computation of clipped/restricted Voronoi diagram

### v2

@@ III-B Edge/face flags of simplex-cell intersection
Our task is to compute a set of simplex-cell intersections as fast as possible. This requires an efficient search for pairs of simplices and Voronoi cells that intersect with each other. Such simplex-cell pairs can be easily found by a propagation approach [46, 47].
@@
The intermediate and final results of each simplex-cell intersection are both a polygon/polytope, and we can attach a flag to each edge/face of the polygon/polytope to achieve the propagation. For an edge/face of polygon/polytope $\mathcal{P}$, if it is shared by $\mathcal{P}$ and a domain simplex $\mathcal{T}_{b}$, its flag is set as $(-1)\times$ the index of $\mathcal{T}_{b}$, i.e., b-b, while if it is shared by $\mathcal{P}$ and a Voronoi cell $\Omega_{j}$, its flag is set as the index of the cell, i.e., $j$. Thus, the edge/face flags of each simplex $\mathcal{T}_{a}\in\mathcal{M}$ can be initialized by using the indices of the adjacent simplices, see Fig.2(a). If $\mathcal{T}_{a}$ intersects with a Voronoi cell $\Omega_{i}$, the intersection $\mathcal{P}_{a,i}$ can be generated after several clippings. Each valid clipping will produce a new edge/face, its flag is set as the index of the neighboring site related to the clipping, see Fig.2(b). Once the intersection $\mathcal{P}_{a,i}$ is finally generated, its edge/face flags indicate either the neighboring Voronoi cells of $\Omega_{i}$ intersected with $\mathcal{T}_{a}$, or the adjacent simplices of $\mathcal{T}_{a}$ intersected with $\Omega_{i}$, so that other simplex-cell pairs can be obtained quickly using edge/face flags of $\mathcal{P}_{a,i}$.
Figure 2: Each edge/face flag indicates the adjacent domain simplex (<0<0) or the neighboring Voronoi cell (>0>0). (a) edge/face flags of a domain simplex $\mathcal{T}_{a}\in\mathcal{M}$; (b) edge/face flags of the intersection $\mathcal{P}_{a,i}$ (light green) of $\mathcal{T}_{a}$ and Voronoi cell $\Omega_{i}$, where the large disks are sites (red, purple, grey), and the small black disks are corner points of $\mathcal{P}_{a,i}$.
(a) $\mathbf{q}\notin\Omega_{i}$
(b) $\mathbf{q}\in\Omega_{i}$
(c) $\mathbf{q}\in\Omega_{i}$
Figure 3: In our method, the point-in-cell test only needs to examine two types of points: the corner points of a polygon/polytope (a–b) and the intersecting points between a bisector and an edge (c). Each blue arrow shows the nearest site $\mathbf{v}$ of a point $\mathbf{q}$.
@@
Figure 4: Illustration of the edge-based bisector search, i.e., find a target bisector on edge $\mathbf{p}_{s}\mathbf{p}_{t}$ from a polygon/polytope. Initially, $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\in\Omega_{j}$ (a), we iteratively compute the intersecting point $\mathbf{q}$ (yellow) of $\mathbf{p}_{s}\mathbf{p}_{t}$ (pink) and the bisector $B_{i,j}$ (purple), and update $j$ as the index of the nearest site of $\mathbf{q}$ (b-c), until $\mathbf{q}\in\Omega_{i}$, then the target bisector is found (d). Using the target bisector to clip the polygon/polytope, the two endpoints of the rest of the edge are both in $\Omega_{i}$ (e). Each blue arrow shows the nearest site $\mathbf{v}$ of a point $\mathbf{p}$. $i=1$ in this figure.
@@
Figure 5: Illustration of computing the intersection $\mathcal{P}$ of a simplex $\mathcal{T}\in\mathcal{M}$ and a Voronoi cell $\Omega_{i}$. At least one corner point of $\mathcal{T}$ should be in $\Omega_{i}$. $\mathcal{P}$ is first initialized as $\mathcal{T}$. Select an edge whose starting point is in $\Omega_{i}$ but the end point is not, then find a bisector by using the proposed edge-based search scheme to clip $\mathcal{P}$. Repeat the selection and clipping steps until all the corner points are in $\Omega_{i}$. $i=1$ in this figure.
@@
Figure 6: Illustration of computing the intersection $\mathcal{P}_{a,j}$ of a simplex $\mathcal{T}_{a}\in\mathcal{M}$ and a Voronoi cell $\Omega_{j}$, where there is no corner point of $\mathcal{T}_{a}$ being in $\Omega_{j}$. (a) The intersection $\mathcal{P}_{a,u}$ (light blue polygon) of $\mathcal{T}_{a}$ and $\Omega_{u}$ has been obtained, one of its edges/faces (red segment) involves a clipping related to the bisector $B_{u,j}$, and its adjacent edges/faces (blue segments) are useful to construct an intermediate polygon/polytope $\mathcal{G}_{a,j}$; (b-c) construction of $\mathcal{G}_{a,j}$; (d) computation of $\mathcal{P}_{a,j}$ from $\mathcal{G}_{a,j}$ by using Algorithm 1. $u=1,j=2$ in this figure.
IV The Proposed Method
Benefiting from the convexity of Voronoi cells, it is known that a polygon/polytope lies entirely in a Voronoi cell if all its corner points are in the cell. If a simplex $\mathcal{T}_{a}\in\mathcal{M}$ intersects with a Voronoi cell $\Omega_{i}$, our strategy is to remove the part of $\mathcal{T}_{a}$ outside $\Omega_{i}$, ensuring all the corner points of the rest are in $\Omega_{i}$.
Each simplex can be decomposed into smaller polygons/polytopes, each belonging to a Voronoi cell, the details are provided in IV-D. Because all Voronoi cells are unknown, how to determine whether a space point is in a target Voronoi cell is the key to our method. We set up a test called point-in-cell in IV-A to accomplish the determination. By using the tests, we present a novel solution that only involves valid clippings to generate a simplex-cell intersection in IV-C, and the bisector of each clipping can be found by an edge-based search scheme introduced in IV-B.
IV-A The point-in-cell test
Given a query point $\mathbf{q}$ and the index $i$ of a Voronoi cell $\Omega_{i}$, the point-in-cell test tells us whether the point is in the cell or not, which is frequently invoked in our method.
Based on the definition of Voronoi cells, $\mathbf{q}$ is in $\Omega_{i}$ if the distance from $\mathbf{q}$ to $\mathbf{v}_{i}$ is not greater than that from $\mathbf{q}$ to any other sites. Let $\mathcal{N}_{\mathbf{q}}$ be the nearest site of $\mathbf{q}$. We can conclude that $\mathbf{q}\in\Omega_{i}$ if $\mathcal{N}_{\mathbf{q}}=\mathbf{v}_{i}$ or if the following equation holds:
$\|\mathbf{q}-\mathbf{v}_{i}\|=\|\mathbf{q}-\mathcal{N}_{\mathbf{q}}\|.$ (2)
Thus, only one time of nearest neighbor search is required in each point-in-cell test.
However, due to the finite precision of floating-point arithmetic in computers, directly applying Equation (2) is unreliable when $\mathcal{N}_{\mathbf{q}}\neq\mathbf{v}_{i}$. To ensure robustness, we therefore perform this judgment using integer flags around $\mathbf{q}$. In our solution for computing the simplex-cell intersection $\mathcal{P}$ of a domain simplex $\mathcal{T}$ and $\Omega_{i}$ (Section IV-C), the point-in-cell test only needs to examine two types of points: (1) the corner points of $\mathcal{P}$ (see Fig. 3(a-b)); (2) the intersecting points between a bisector and an edge of $\mathcal{P}$ (see Fig. 3(c)).
For a corner point $\mathbf{q}$ of $\mathcal{P}$, we collect all flags $\mathcal{F}_{\mathbf{q}}$ of edges/faces incident to $\mathbf{q}$, and if the index of $\mathcal{N}_{\mathbf{q}}$ appears in $\mathcal{F}_{\mathbf{q}}$, then $\mathbf{q}\in\Omega_{i}$ for the case $\mathcal{N}_{\mathbf{q}}\neq\mathbf{v}_{i}$. For example in Fig. 3(a), $\mathcal{F}_{\mathbf{q}}=\left\{-c,k\right\}$, $\mathcal{N}_{\mathbf{q}}=\mathbf{v}_{j}$, because $j\notin\mathcal{F}_{\mathbf{q}}$, then $\mathbf{q}\notin\Omega_{i}$; while in Fig. 3(b), $\mathcal{F}_{\mathbf{q}}=\left\{j,k\right\}$, $\mathcal{N}_{\mathbf{q}}=\mathbf{v}_{j}$, we have $\mathbf{q}\in\Omega_{i}$ since $j\in\mathcal{F}_{\mathbf{q}}$.
For an intersecting point $\mathbf{q}$ between a bisector $B_{i,j}$ of sites $\mathbf{v}_{i}$, $\mathbf{v}_{j}$ and an edge, if $\mathcal{N}_{\mathbf{q}}=\mathbf{v}_{j}$ (Fig. 3(c)), it is known that $\mathbf{q}$ must be on the common edge of Voronoi cells $\Omega_{i}$ and $\Omega_{j}$, namely, $\mathbf{q}\in\Omega_{i}$.
Algorithm 1 Computation of simplex-cell intersection
0: a simplex $\mathcal{T}\in\mathcal{M}$ and a site index ii.
@@ IV-A The point-in-cell test
3:  select an edge $\mathbf{p}_{s}\mathbf{p}_{t}$, requiring that $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\notin\Omega_{i}$;
4:  find a bisector $B_{i,j}$ on $\mathbf{p}_{s}\mathbf{p}_{t}$ (Section IV-B);
@@
IV-B Edge-based bisector search
It is known that the points shared by two neighboring Voronoi cells $\Omega_{u},\Omega_{w}$ must be on the bisector $B_{u,w}$ of the two corresponding sites $\mathbf{v}_{u},\mathbf{v}_{w}$, and the clipping using $B_{u,w}$ must contribute to the computation of Voronoi cells $\Omega_{u},\Omega_{w}$. This observation inspires us the way to search for the target bisector on edge $\mathbf{p}_{s}\mathbf{p}_{t}$.
Initially, $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\in\Omega_{j}$, we construct the bisector $B_{i,j}$ of the corresponding sites $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, and compute the intersecting point $\mathbf{q}$ of $\mathbf{p}_{s}\mathbf{p}_{t}$ and $B_{i,j}$. By using the point-in-cell test (section IV-A), we can easily obtain the relationship between $\mathbf{q}$ and $\Omega_{i}$: (1) If the test answers false, and the nearest site of $\mathbf{q}$ is denoted as $\mathbf{v}_{k}$, then $k$ must be different with $i$ and $j$, which means that the clipping using $B_{i,j}$ must be useless. In this case, we update $j\leftarrow k,\mathbf{p}_{t}\leftarrow\mathbf{q}$ and repeat the above steps, i.e., update $B_{i,j}$ and $\mathbf{q}$, and redo the point-in-cell test. (2) If the test answers true, we know that the whole segment $\mathbf{p}_{s}\mathbf{q}$ is in the cell $\Omega_{i}$, then the newest $B_{i,j}$ is the target bisector, which can be directly used to clip the polygon/polytope. Fig. 4 shows an illustration.
Algorithm 2 Computation of clipped or restricted VD
@@ IV-B Edge-based bisector search
1: build a structure $\mathcal{S}$ of $\mathcal{V}$ for nearest neighbor searches (section VI-B);
@@
3:for each $\mathcal{T}_{a}\in\mathcal{M}$ do
@@
4:  $\mathbf{v}_{i}\leftarrow$ the nearest site of the first corner point of $\mathcal{T}_{a}$;
5:  compute intersection $\mathcal{P}_{a,i}$ of $\mathcal{T}_{a}$ and $\Omega_{i}$ (Alg. 1);
6:  $\Omega_{i|\mathcal{M}}\leftarrow\Omega_{i|\mathcal{M}}\bigcup\mathcal{P}_{a,i}$;
7:  $Q\leftarrow\emptyset$;
8:  add other sites to $Q$ using edge/face flags of $\mathcal{P}_{a,i}$;
9:  while $Q$ is not empty do
10:   $\mathbf{v}_{j}\leftarrow$ popped from $Q$;
11:   construct a polygon/polytope $\mathcal{G}_{a,j}$ from $\mathcal{T}_{a}$;
12:   compute intersection $\mathcal{P}_{a,j}$ of $\mathcal{G}_{a,j}$ and $\Omega_{j}$ (Alg. 1);
13:   $\Omega_{j|\mathcal{M}}\leftarrow\Omega_{j|\mathcal{M}}\bigcup\mathcal{P}_{a,j}$;
14:   add unvisited sites to $Q$ using edge/face flags of $\mathcal{P}_{a,j}$;
15:  end while
16:end for
The target bisector on the edge can be found after several point-in-cell tests. The number of tests is proportional to the number of Voronoi cells intersecting the edge. To reduce the number of tests, a minor modification can be made before searching for the target bisector, which will be discussed later in Section VI-A.
IV-C Computation of simplex-cell intersection
Suppose that at least one corner point of a simplex $\mathcal{T}$ is in the Voronoi cell $\Omega_{i}$ of site $\mathbf{v}_{i}$, we present a novel solution for generating the intersection $\mathcal{P}$ of $\mathcal{T}$ and $\Omega_{i}$. Specifically, $\mathcal{P}$ is initialized as $\mathcal{T}$, and it is iteratively removed the part outside $\Omega_{i}$. The steps are as follows: (1) select an edge $\mathbf{p}_{s}\mathbf{p}_{t}$ from $\mathcal{P}$, requiring that $\mathbf{p}_{s}\in\Omega_{i},\mathbf{p}_{t}\notin\Omega_{i}$; (2) find a bisector $B_{i,j}$ by using the proposed edge-based search scheme (Section IV-B) on $\mathbf{p}_{s}\mathbf{p}_{t}$, and use it to clip $\mathcal{P}$, the result is still denoted as $\mathcal{P}$. Repeat the selection and clipping steps until $\mathcal{P}$ has no edge meeting the requirement in step 1, which implies that all the corner points of $\mathcal{P}$ are in $\Omega_{i}$, i.e., the simplex-cell intersection is finally obtained. The steps are also listed in Algorithm 1, and an illustration is given in Fig. 5.
At most one clipping is involved for each edge of the intersection, and all the clippings are contributed to the final result. This is a great advantage of our solution, since a part of the clippings in both DT-based and kNN-based methods could be useless for generating such intersections. Furthermore, when clipping $\mathcal{P}$ with a bisector, a part of corner points of $\mathcal{P}$ are guaranteed to lie in the Voronoi cell $\Omega_{i}$. Thus, it is unnecessary to check whether these corner points belong to the half-space containing $\mathbf{v}_{i}$, which reduces computational effort.
IV-D Computation of clipped or restricted Voronoi diagram
We now concentrate on the computation of clipped or restricted Voronoi cells $\left\{\Omega_{i|\mathcal{M}}\right\}_{i=1}^{n}$. Initially, they are set to empty, i.e., $\Omega_{i|\mathcal{M}}=\emptyset,i=1,\dots,n$. We can decompose each domain simplex into smaller polygons/polytopes, each of which is entirely in a Voronoi cell such that it can be added to the corresponding clipped or restricted cell. Since the decomposition process is identical for all simplices, we take $\mathcal{T}_{a}\in\mathcal{M}$ as an example to describe our method. The pseudo-code can be found in Algorithm 2.